# AlignDiT: Multimodal Aligned Diffusion Transformer for Synchronized Speech Generation

### v1

@@ 3.3 Training Objective
Second, to ensure that AlignDiT not only generates natural mel-spectrograms but also learns effective modality alignment, relying solely on the CFM loss can be insufficient, as it provides only an indirect learning signal. To address this, we guide the intermediate representations of the DiT blocks to align more directly with the textual content. Specifically, we introduce a Connectionist Temporal Classification (CTC) [21] loss by attaching a lightweight projection head to an intermediate DiT block, which predicts the text sequence from its hidden representations. It can be expressed as:
$\mathcal{L}_{\text{CTC}}=-\log p_{\text{CTC}}(text\,|\,h^{i}),$ (8)
where $h^{i}$ is the output of $i$-th DiT block. The CTC loss encourages the model to retain more linguistic information, without employing external alignment modules.
The total multi-task loss is defined as follows:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CFM}}+\lambda_{\text{CTC}}\mathcal{L}_{\text{CTC}},$ (9)
@@
3.4 Audio-only Pretraining
Generating high-quality speech and learning multimodal alignment simultaneously can be challenging. To better initialize AlignDiT model, we employ an audio-only pretraining phase before conditioning video and text modalities. The model learns to predict masked regions of the input mel-spectrogram using the unmasked regions during pretraining and this has demonstrated its effectiveness in text‐to‐speech task [42]. In addition, we follow the approach [76] that improves convergence speed by distilling rich features into the intermediate layers of DiT, using a self-supervised speech model [29] as the teacher. By allowing the model to acquire unconditional speech generation abilities prior to multimodal conditioning, this pretraining phase simplifies the subsequent audio-video-text fusion. Additionally, since audio‐only data is abundant compared to audio-video-text paired data, our audio‐only pretraining strategy is readily scalable and offers a practical way to improve the performance.
3.5 Multimodal Classifier-Free Guidance
In diffusion-based generative models, classifier-free guidance [25] is well-explored to strengthen the influence of conditioning signals during inference. This is achieved by using both conditional and unconditional predictions from the same model to guide the generation process as follows:
$v_{t,CFG}=v_{t}(x_{t},h)+s\cdot(v_{t}(x_{t},h)-v_{t}(x_{t},\varnothing)),$ (10)
@@ 3.5 Multimodal Classifier-Free Guidance
$\begin{split}v_{t,CFG}=&\>v_{t}(x_{t},h)\\ &+s_{t}\cdot(v_{t}(x_{t},h_{text})-v_{t}(x_{t},\varnothing))\\ &+s_{v}\cdot(v_{t}(x_{t},h)-v_{t}(x_{t},h_{text})),\end{split}$ (11)
where $s_{t}$ is guidance scale for text modality and $s_{v}$ for remained modality, namely video. By adjusting $s_{t}$ and $s_{v}$, we can adaptively control the focus between modalities. Higher $s_{t}$ encourages the model to follow the text more closely, improving intelligibility, while higher $s_{v}$ leads to better lip synchronizations.
@@ 4.1 Datasets
We train and test our proposed AlignDiT on the large-scale audio-visual LRS3 dataset [2]. LRS3 is an English sentence-level dataset sourced from TED and TEDx talk videos. Each video in LRS3 is accompanied by an accurate transcript. It comprises approximately 439 hours of videos with unconstrained, long utterances from thousands of speakers, making it suitable for our multimodal setting. Around 131,000 utterances are utilized for training. For testing, we construct an LRS3-cross set consisting of triplets of {reference speech, text, silent video}, where the reference speech is from a different utterance by the same speaker, rather than using the ground-truth speech as the reference. This prevents the model from accessing ground-truth speech during testing, ensuring a rigorous evaluation without information leakage. For video-to-speech, instead of using the provided text from the dataset, we utilize an off-the-shelf lip reading model to extract transcripts directly from the silent videos.
4.2 Evaluation Metrics
@@
Subjective metrics. Since our primary focus is generating natural and synchronized speech, subjective evaluation is essential for accurately assessing model performance [47, 3]. Thus, we conduct human evaluations using Mean Opinion Scores (MOS) for the ADR task along two criteria: naturalness, evaluating the overall quality of the speech; and similarity, assessing speaker similarity between the reference speech and synthesized speech. 20 participants rate the randomly sampled 30 utterances using a 5-point Likert scale, where 1 indicates “very poor” (or “very different”) and 5 indicates “very good” (or “very similar”).
@@
Objective metrics. We employ comprehensive objective metrics to evaluate the synthesized speech. For both ADR and video-to-speech tasks, we report Word Error Rate (WER) to assess content accuracy, using Whisper-large-v3 [61] to transcribe the synthesized speech. Speaker similarity (spkSIM) is measured by computing the cosine similarity between speaker embeddings extracted from synthesized and reference speech using a WavLM-large-based speaker verification model [6]. Lip-sync accuracy (AVSync) between the video and synthesized speech is measured using AVHubert [65]. Specifically, we compute the cosine similarity between AVHubert features extracted from video paired with ground-truth speech and features extracted from video paired with synthesized speech. This assessment has been shown in [72] to be more robust and effective for validating audio-visual synchronization compared to conventional lip-sync accuracy metrics based on SyncNet [13]. For the alignment task, we evaluate alignment accuracy by comparing word-level timestamps obtained from our method with ground truth timestamps derived from original audio alignment. We report the average absolute time error per word (in milliseconds).
4.3 Implementation Details
Data preprocessing. For audio features, we use 16 kHz mono audio and convert it into 80 bins mel-spectrogram using a filter size of 640, a hop size of 160, resulting in a frame rate of 100 Hz. We utilize 25 fps video and extract visual features as follows. Face detection is performed using RetinaFace [17], followed by facial landmark extraction using FAN [5]. We crop the lip-centered region based on the detected landmarks and resize it into 88 × 88. The pretrained AV-HuBERT (Large) model [65] is utilized to extract 25 Hz visual representations, which result in a fixed 1:4 length ratio to the audio features. For text, we represent input as a sequence of characters. Compared to the audio features, the character sequence generally has a shorter temporal length [62].
Architecture. The visual feature encoder is composed of two transposed convolution layers with stride 2, followed by two Conformer [22] encoder layers, each with embedding size of 512, 4 attention heads, and a 1024-dimensional feed-forward layer. For the text encoder, we adopt 4 layers of ConvNeXt v2 [70] with 512-dimensional hidden embeddings. After concatenating multimodal features, we apply a linear layer to project the fused representation. This representation is then processed by 18 DiT blocks, each with 768-dimensional embedding size, 12 attention heads, and a 3072-dimensional feed-forward layer.
Training. We train AlignDiT using AdamW [43] optimizer with a warmup of 20k steps to a peak learning rate of $7.5\times 10^{-5}$, followed by linearly decay. Pretraining is conducted for 500k steps using audio-only data with a total batch size of 0.3 hours. We then finetune the model for 400k steps using audio, video, and text paired data with a total batch size of 0.1 hours. $\lambda_{CTC}$ is set to 0.1 to balance the CTC loss against the CFM loss in initial stage. The modality dropout probability is set to 0.2 for text, video, and all modalities, respectively.
@@ 4.3 Implementation Details
Table 1: Ablation study of modality conditioning method.
@@
Table 2: Ablation study of audio-only pretraining.
Inference. During inference, we apply Exponential Moving Averaged (EMA) weights with a decay rate of 0.999 to stabilize the model prediction. AlignDiT takes a reference speech along with its corresponding transcript as inputs, and the total duration is determined by the number of video frames (4 times longer than the video frames). For sampling, we use the Euler ODE solver with timestep scaling based on sway sampling strategy with a coefficient of -1, following F5-TTS [8]. To convert generated mel spectrograms into waveforms, we employ a HiFi-GAN [36] model trained on the LRS3 dataset.
4.4 Baseline Models
We compare AlignDiT with state-of-the-art open-source ADR systems, HPMDubbing [14] and StyleDubber [15]. To ensure a fair comparison and improve their generalization capability, we train these models on the LRS3 dataset [2], which is significantly larger than the datasets originally used in their training. Additionally, we replace the lip feature extractor in each model with AV-HuBERT (large) [65], unifying the lip feature extractor across AlignDiT, HPMDubbing, and StyleDubber to purely evaluate and compare their effectiveness under identical settings. Note that both baseline models require an explicit duration aligner during training. For video-to-speech, we compare our method with DiffV2S [9], Intelligible [10], and LipVoicer [74], which are specifically designed for this task. For visual forced alignment, we compare against visual keyword spotting methods such as KWS-Net [53] and Transpotter [60], as well as CTC-based methods [38] and DVFA [35].
$\lambda_{t}$ $\lambda_{v}$ WER $\downarrow$ AVSync $\uparrow$ spkSIM $\uparrow$
@@ 4.4 Baseline Models
Table 3: Ablation study of multimodal CFG.
@@
Table 4: Ablation study of input modalities.
Figure 3: We qualitatively compare mel-spectrogram visualizations of ground-truth speech and synthesized speech from the ADR task using AlignDiT (ours) and existing methods, HPMDubbing and StyleDubber. We also provide results from unimodal inputs, either video-only (VTS) or text-only (TTS). The text below each mel-spectrogram represents time-aligned speech extracted using Whisper-large-v3, with red text indicating incorrectly synthesized word compared to the ground truth.
@@
GT 4.13$\pm$0.13 3.96$\pm$0.14 2.335 1.000 0.562
HPMDubbing [14] 2.37$\pm$0.14 2.43$\pm$0.16 5.382 0.750 0.287
StyleDubber [15] 1.79$\pm$0.12 2.35$\pm$0.18 3.170 0.586 0.349
Ours 3.79$\pm$0.16 3.96$\pm$0.13 1.401 0.751 0.515
Table 5: Experimental results on LRS3-cross set. Subjective MOS results are presented with 95% confidence interval. $\uparrow$ represents that a higher score is better, and $\downarrow$ denotes that a lower score is better. The best-performing result is shown in bold.
5 Results
5.1 Ablation Studies
We first demonstrate the contribution of each AlignDiT component through in-depth ablation studies. We conduct a series of experiments using various objective metrics, i.e. WER, AVSync, and spkSIM, and present comprehensive findings on modality conditioning method, audio-only pretraining, multimodal CFG, and input modalities.
@@ 5.1 Ablation Studies
Audio-only pretraining. Table 2 illustrates the benefits of audio-only pretraining. Pretraining AlignDiT on the LRS3 dataset, which is also used during the main training phase, shows consistent quality improvements across all evaluation metrics. Even greater improvements are achieved when using LibriSpeech [54], a dataset not seen during the subsequent training. These results underscore the benefits of leveraging diverse audio data during pretraining, and demonstrate that the performance of AlignDiT can be readily enhanced using easily accessible audio resources.
Multimodal CFG. In order to explore how AlignDiT benefits from its multimodal CFG, we experiment with different values of $s_{t}$ and $s_{v}$. From the results in Table 3, we derive two key findings. First, applying CFG ($s_{t}=s_{v}\in\{2,5\}$) consistently improves performance compared to not using CFG ($s_{t}=s_{v}=0$), aligning with observations from prior studies [25, 39]. Second, and more importantly, leveraging the proposed multimodal CFG, i.e., adjusting $s_{t}$ and $s_{v}$ to balance information from each modality, further enhances overall performance, demonstrating its effectiveness in synthesizing more natural speech. Through analysis of optimal guidance scales for each modality, we set $s_{t}$ = 5 and $s_{t}$ = 2, which achieves the best quality in terms of WER and spkSIM.
@@ 5.2 Quantitative Comparison
DiffV2S [9] 35.210 0.608 0.115
Intelligible [10] 27.432 0.675 0.316
LipVoicer [74] 21.164 0.524 0.094
@@
Table 6: Video-to-Speech benchmark.
5.3 Qualitative Comparison
We visually compare the mel-spectrograms converted from synthesized speech produced by existing ADR methods [14, 15] and our AlignDiT with multimodal input, alongside those from ground-truth speech, in Fig. 3. We also include results synthesized from unimodal inputs: video-only (VTS) and text-only (TTS). Focusing on the orange and red boxes in columns 1 and 2, as well as the transcribed speech below each mel-spectrogram, we observe that existing methods frequently generate incorrect speech and blurry spectrograms. In columns 3 and 4, although existing methods produce accurate speech content, they occasionally generate unintended sounds or produce mel-spectrograms that lack details and appear blurry. Regarding VTS and TTS, we observe consistent results in Sec. 5.1. For VTS across all samples, since no text input is provided, content accuracy becomes significantly unstable, leading to incorrect speech. Conversely, TTS generates accurate speech content across all examples due to the text input, yet the absence of video input results in temporal misalignment compared to the ground-truth. In contrast, our model consistently synthesizes accurate speech content, exhibiting fine details in the mel-spectrogram that closely match the ground-truth across all examples. These results clearly demonstrate that our model effectively synthesizes high-quality, accurate speech comparable to the ground-truth.
5.4 Applications

### v2

@@ 3.3. Training Objective
Second, while the CFM loss encourages natural mel-spectrogram generation, relying solely on the CFM loss can be insufficient for modality alignment, as it provides only an indirect learning signal (Choi et al., 2025b). To address this, we introduce a Connectionist Temporal Classification (CTC) (Graves et al., 2006) loss to guide the intermediate representations of DiT blocks to align more directly with the textual content. Specifically, we attach lightweight projection heads to several intermediate blocks to predict the text sequence from their hidden representations. The loss can be expressed as:
(8) $\mathcal{L}_{\text{CTC}}=-\sum_{i\in\mathcal{I}}\log p_{\text{CTC}}(\text{text}\mid h^{i}),$
where $h^{i}$ is the output of the $i$-th DiT block and $\mathcal{I}$ denotes the selected layers. The CTC loss encourages the model to retain more linguistic information, without employing external alignment modules. The total multi-task loss is defined as follows:
(9) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CFM}}+\lambda_{\text{CTC}}\mathcal{L}_{\text{CTC}},$
@@
3.4. Audio-only Pretraining
Generating high-quality speech and learning multimodal alignment simultaneously can be challenging. To provide a better initialization for AlignDiT, we employ an audio-only pretraining phase before conditioning on video and text. The model learns to predict masked regions of the input mel-spectrogram from unmasked context during pretraining, and this has demonstrated its effectiveness in text‐to‐speech task (Liu et al., 2024). In addition, we follow the approach (Yu et al., 2025) that improves convergence speed by distilling rich features into the intermediate layers of DiT, using a self-supervised speech model (Hsu et al., 2021) as the teacher. By allowing the model to acquire unconditional speech generation abilities prior to multimodal conditioning, this pretraining phase simplifies the subsequent audio-video-text fusion. Additionally, since audio‐only data is abundant compared to audio-video-text paired data, our audio‐only pretraining strategy is readily scalable and offers a practical way to improve the performance.
3.5. Multimodal Classifier-Free Guidance
In diffusion-based generative models, classifier-free guidance (Ho and Salimans, 2021) is well-explored to strengthen the influence of conditioning signals during inference. This is achieved by using both conditional and unconditional predictions from the same model to guide the generation process as follows:
(10) $v_{t,CFG}=v_{t}(x_{t},h)+s\cdot(v_{t}(x_{t},h)-v_{t}(x_{t},\emptyset)),$
@@ 3.5. Multimodal Classifier-Free Guidance
(11) $\begin{split}v_{t,CFG}=&\>v_{t}(x_{t},h_{text},h_{video})\\ &+s_{video}\cdot(v_{t}(x_{t},h_{text},h_{video})-v_{t}(x_{t},h_{text},\emptyset))\\ &+s_{text}\cdot(v_{t}(x_{t},h_{text},\emptyset)-v_{t}(x_{t},\emptyset,\emptyset)),\end{split}$
where $s_{text}$ is guidance scale for text modality and $s_{video}$ for video. By adjusting $s_{text}$ and $s_{video}$, we can adaptively control the focus between modalities. Higher $s_{text}$ encourages the model to follow the text more closely, improving intelligibility, while higher $s_{video}$ leads to better lip synchronizations.
@@ 4.1. Datasets
We train and evaluate our proposed AlignDiT on the large-scale audio-visual dataset LRS3 (Afouras et al., 2018a), which contains 439 hours of English sentence-level data sourced from TED and TEDx talk videos from thousands of speakers. Each video segment contains unconstrained audio-visual speech paired with an accurate transcript, making the dataset suitable for our multimodal setting. Around 131,000 utterances are utilized for training. For evaluation, we construct an LRS3-cross test set consisting of triplets of {reference speech, text, silent video}, where the reference speech is from a different utterance by the same speaker, rather than using the ground-truth speech as the reference. This prevents the model from accessing ground-truth speech during testing, ensuring a rigorous evaluation without information leakage. For video-to-speech, instead of using the provided text from the dataset, we utilize an off-the-shelf lip reading model to extract transcripts directly from the silent videos.
4.2. Evaluation Metrics
@@
Subjective metrics. Since our primary focus is generating natural and synchronized speech, subjective evaluation is essential for accurately assessing model performance (Maiti et al., 2023; Ahn et al., 2024). Thus, we conduct human evaluations using Mean Opinion Scores (MOS) for the ADR task along two criteria: naturalness, evaluating the overall quality of the speech; and similarity, assessing speaker similarity between the reference speech and synthesized speech. 20 participants rate the randomly sampled 30 utterances using a 5-point Likert scale, where 1 indicates “very poor” (or “very different”) and 5 indicates “very good” (or “very similar”).
@@
Objective metrics. We employ comprehensive objective metrics to evaluate the synthesized speech. For both ADR and video-to-speech tasks, we report Word Error Rate (WER) to assess content accuracy, using Whisper-large-v3 (Radford et al., 2023) to transcribe the synthesized speech. Speaker similarity (spkSIM) is measured by computing the cosine similarity between speaker embeddings extracted from synthesized and reference speech using a WavLM-large-based speaker verification model (Chen et al., 2022). Lip-sync accuracy (AVSync) between the video and synthesized speech is measured using AVHubert (Shi et al., 2022). Specifically, we compute the cosine similarity between AVHubert features extracted from video paired with ground-truth speech and features extracted from video paired with synthesized speech. This assessment has been shown in (Yaman et al., 2024) to be more robust and effective for validating audio-visual synchronization compared to conventional lip-sync accuracy metrics based on SyncNet (Chung and Zisserman, 2016). For the alignment task, we evaluate alignment accuracy by comparing word-level timestamps obtained from our method with ground truth timestamps derived from original audio alignment. We report the average absolute time error per word (in milliseconds).
Table 1. Ablation study of modality conditioning method.
@@
Table 2. Ablation study of audio-only pretraining.
@@
4.3. Implementation Details
Data preprocessing. For audio features, we use 16 kHz mono audio and convert it into 80 bins mel-spectrogram using a filter size of 640, a hop size of 160, resulting in a frame rate of 100 Hz. We utilize 25 fps video and extract visual features as follows. Face detection is performed using RetinaFace (Deng et al., 2020), followed by facial landmark extraction using FAN (Bulat and Tzimiropoulos, 2017). We crop the lip-centered region based on the detected landmarks and resize it into 88 × 88. The pretrained AV-HuBERT (Large) model (Shi et al., 2022) is utilized to extract 25 Hz visual representations, which result in a fixed 1:4 length ratio to the audio features. For text, we represent input as a sequence of characters. Compared to the audio features, the character sequence generally has a shorter temporal length (Ren et al., 2019).
Architecture. The visual feature encoder is composed of two transposed convolution layers with stride 2, followed by two Conformer (Gulati et al., 2020) encoder layers, each with embedding size of 512, 4 attention heads, and a 1024-dimensional feed-forward layer. For the text encoder, we adopt 4 layers of ConvNeXt v2 (Woo et al., 2023) with 512-dimensional hidden embeddings. After concatenating multimodal features, we apply a linear layer to project the fused representation. This representation is then processed by 18 DiT blocks, each with 768-dimensional embedding size, 12 attention heads, and a 3072-dimensional feed-forward layer.
Training. We train AlignDiT using AdamW (Loshchilov and Hutter, 2019) optimizer with a warmup of 20k steps to a peak learning rate of $7.5\times 10^{-5}$, followed by linearly decay. Pretraining is conducted for 500k steps on audio-only data with a total batch size of 0.3 hours. We finetune the model for 400k steps using paired audio, video, and text data with a total batch size of 0.1 hours. $\lambda_{CTC}$ is set to 0.1 to balance the CTC loss against the CFM loss in initial stage. The modality dropout probability is set to 0.2 for text, video, and all modalities, respectively.
Inference. During inference, we apply Exponential Moving Averaged (EMA) weights with a decay rate of 0.999 to stabilize the model prediction. AlignDiT takes a reference speech along with its corresponding transcript as inputs, and the total duration is determined by the input video length. For sampling, we use the Euler ODE solver with timestep scaling based on sway sampling strategy with a coefficient of -1, following F5-TTS (Chen et al., 2025). To convert generated mel spectrograms into waveforms, we employ a HiFi-GAN (Kong et al., 2020) model trained on the LRS3 dataset.
4.4. Baseline Models
We compare AlignDiT with state-of-the-art open-source ADR systems, HPMDubbing (Cong et al., 2023) and StyleDubber (Cong et al., 2024). To ensure a fair comparison and improve their generalization capability, we train these models on the LRS3 dataset (Afouras et al., 2018a), which is significantly larger than the datasets originally used in their training. Additionally, we replace the lip feature extractor in each model with AV-HuBERT (large) (Shi et al., 2022), unifying the lip feature extractor across AlignDiT, HPMDubbing, and StyleDubber to purely evaluate and compare their effectiveness under identical settings. Note that both baseline models require an explicit duration aligner during training. For video-to-speech, we compare our method with DiffV2S (Choi et al., 2023a), Intelligible (Choi et al., 2023b), and LipVoicer (Yemini et al., 2024), which are specifically designed for this task. For visual forced alignment, we compare against visual keyword spotting methods such as KWS-Net (Momeni et al., 2020) and Transpotter (Prajwal et al., 2021), as well as CTC-based methods (Kürzinger et al., 2020) and DVFA (Kim et al., 2023).
Table 3. Ablation study of multimodal CFG.
$s_{text}$ $s_{video}$ WER $\downarrow$ AVSync $\uparrow$ spkSIM $\uparrow$
@@ 4.4. Baseline Models
Table 4. Ablation study of input modalities.
@@
Figure 3. We qualitatively compare mel-spectrogram visualizations of ground-truth speech and synthesized speech from the ADR task using AlignDiT (ours) and existing methods, HPMDubbing and StyleDubber. We also provide results from unimodal inputs, either video-only (VTS) or text-only (TTS). The text below each mel-spectrogram represents time-aligned speech extracted using Whisper-large-v3, with red text indicating incorrectly synthesized word compared to the ground truth.
Table 5. Experimental results on LRS3-cross set. Subjective MOS results are presented with 95% confidence interval. $\uparrow$ represents that a higher score is better, and $\downarrow$ denotes that a lower score is better. The best-performing result is shown in bold.
@@
[…]
5. Results
5.1. Ablation Studies
We analyze the contribution of each AlignDiT component through in-depth ablation studies. We conduct a series of experiments using various objective metrics, i.e. WER, AVSync, and spkSIM, and present comprehensive findings on modality conditioning method, audio-only pretraining, multimodal CFG, and input modalities.
@@ 5.1. Ablation Studies
Audio-only pretraining. Table 2 illustrates the benefits of audio-only pretraining. Pretraining AlignDiT on the LRS3 dataset, which is also used during the main training phase, shows consistent quality improvements across all evaluation metrics. Even greater improvements are achieved when using LibriSpeech (Panayotov et al., 2015), a dataset not seen during the subsequent training. These results underscore the benefits of leveraging diverse audio data during pretraining, and demonstrate that the performance of AlignDiT can be readily enhanced using easily accessible audio resources.
Multimodal CFG. In order to explore how AlignDiT benefits from its multimodal CFG, we experiment with different values of $s_{text}$ and $s_{video}$. From the results in Table 3, we derive two key findings. First, applying CFG ($s_{text}=s_{video}\in\{2,5\}$) consistently improves performance compared to not using CFG ($s_{text}=s_{video}=0$), aligning with observations from previous works (Ho and Salimans, 2021; Le et al., 2023). More importantly, leveraging the proposed multimodal CFG, i.e., adjusting $s_{text}$ and $s_{video}$ to balance information from each modality, further enhances overall performance, demonstrating its effectiveness in synthesizing more natural speech. Through analysis of optimal guidance scales for each modality, we set $s_{text}$ = 5 and $s_{text}$ = 2, which achieves the best quality in terms of WER and spkSIM.
@@ 5.2. Quantitative Comparison
DiffV2S (Choi et al., 2023a) 35.210 0.608 0.115
Intelligible (Choi et al., 2023b) 27.432 0.675 0.316
LipVoicer (Yemini et al., 2024) 21.164 0.524 0.094
@@
5.3. Qualitative Comparison
We visually compare the mel-spectrograms converted from synthesized speech produced by existing ADR methods (Cong et al., 2023, 2024) and our AlignDiT with multimodal input, alongside those from ground-truth speech, in Fig. 3. We also include results synthesized from unimodal inputs: video-only (VTS) and text-only (TTS). Focusing on the orange and red boxes in columns 1 and 2, as well as the transcribed speech below each mel-spectrogram, we observe that existing methods frequently generate incorrect speech and blurry spectrograms. In columns 3 and 4, although existing methods produce accurate speech content, they occasionally generate unintended sounds or produce mel-spectrograms that lack details and appear blurry. Regarding VTS and TTS, we observe consistent results in Sec. 5.1. For VTS across all samples, since no text input is provided, content accuracy becomes significantly unstable, leading to incorrect speech. Conversely, TTS generates accurate speech content across all examples due to the text input, yet the absence of video input results in temporal misalignment compared to the ground-truth. In contrast, our model consistently synthesizes accurate speech content, exhibiting fine details in the mel-spectrogram that closely match the ground-truth across all examples. These results clearly demonstrate that our model effectively synthesizes high-quality, accurate speech comparable to the ground-truth.
5.4. Applications